# Seeded Poisson Factorization: Leveraging domain knowledge to fit topic models

### v1

@@ 3.2 Post-processing and final inference
The final topic assignment is based on a Naive Bayes classifier (see, e.g., Zhang,, 2004), i.e., the document is assigned to the topic where the per-document posterior mean estimate is maximum.
@@ 3.3 Computational details
The SPF model is implemented in Python 3.10. It allows Graphics Processing Unit (GPU) support due to its implementation in the TensorFlow environment. The model’s source code is mainly based on TensorFlow 2.15 as well as TensorFlow’s add-on library for probabilistic reasoning, TensorFlow Probability 0.23.0. In its standard implementation, SPF uses a batch size of 1 02410241\,0241 024 documents and trains the model for $150$ epochs. Leveraging TensorFlow’s computational graph and gradient tape functionalities, the implementation enables efficient tracking of operations for automatic differentiation during model training.
@@ 4 Empirical results
4.1 Illustrative application
[…]
When fitting the SPF model, we set the number of topics to correspond to the number of product categories, i.e., $K=6$, and follow the lines of Watanabe and Zhou, (2022) to construct a balanced lexicon of ten frequently occurring seed words for each topic. To generate these seed words, we computed the TF-IDF (Term Frequency-Inverse Document Frequency) values for each word within each category. We select the top 10 words from the TF-IDF matrix as the seed words for each topic. This process is conducted automatically to minimize subjectivity and to ensure that the seed word selection process remains as objective as possible. Table 1 presents a summary of the final sample of documents and the pre-specified seed words per product category. We use the standard specifications for prior and variational parameter values as outlined in Sections 2 and 3 and set the learning rate to $0.1$, the number of epochs to $150$ and the batch size to 1 02410241\,0241 024.
@@ 4.1 Illustrative application
4.1.1 Topic interpretability
We examine the interpretability of the topics inferred by SPF based on the approximate posterior mean topic-term intensities $\hat{\bm{\beta}}_{k}$ for each topic $k$. Table 2 presents the top 14 terms with the highest approximate mean intensities per topic, after removing stopword-like terms that provide no contextual information. Bold terms represent seed words. The mean term intensity (provided in parentheses) is calculated as the sum of the seeded term intensity and the unseeded term intensity, as defined in Equation 7. These high-intensity words per topic enable the characterization of the topic as well as assess how influential the seed words were.
Clearly the pre-defined seed words exhibit a strong presence among the most pertinent terms for all topics. However, the number of seed words included in the top 14 words with highest intensity varies across topics. For topic ‘Toys‘, all 10 seed words are also included in the list of 14 most pertinent terms for this topic. This number decreases to six for ‘Health’, eight for ‘Beauty’, seven for ‘Baby’ and eight for ‘Pets’. The lowest number of seed words are included in the list of 14 most pertinent terms for the topic ‘Grocery’ where only four out of the ten seed words are listed.
Table 2 reveals that the model effectively prioritizes not only the explicitly defined seed words but also identifies and assigns significant weight to relevant additional terms that are not pre-specified as seed words. E.g., for topic ‘Health’ the seed word ‘day’ was specified and also ‘days’ is included among the 14 most pertinent terms. For topic ‘Baby’ not only the seed word ‘son’ is included but also ‘daughter’. Inspecting the ‘Baby’ topic further by also assessing additional terms with high intensity indicates that SPF did not only assign high relevance to expected seed words such as ‘seat’ (14.88) and ‘son’ (12.07), but also recognized terms like ‘bed’ (6.71) and ‘sleep’ (6.31) as highly pertinent to the topic. These terms align with the product subcategory ‘sleep positioners’, which falls under the broader ‘Baby’ category, demonstrating the model’s nuanced understanding of topic content and its ability to discern contextually important terms.
The topic ‘Grocery’ fails to capture most of the seed words among the 14 most pertinent terms. However, the terms with high intensity suggest that this topic captured an additional aspect in customer reviews which relates not to the product but the purchase process. For example, terms like ‘store’ ($7.34$) and ‘shipping’ ($7.20$) have a high prevalence in the ‘Grocery’ topic.
@@ 4.1.1 Topic interpretability
Table 2 also illustrates that SPF is able to assign distinctly different intensities to the seed words as well as other terms with high intensity within their respective topics. E.g., within the ‘Pets’ category, ‘dog’ ($24.57$) plays a dominant role, whereas other seed words, like ‘food’ ($7.41$), display a markedly lower mean intensity. This contrast highlights the ability of SPF to estimate the uneven influence of seed words in defining a topic. This property of the SPF topic model is important to also mitigate the risk of potential mis-specifications that may arise due to incomplete domain knowledge during the initial selection of seed words. To empirically assess the influence of mis-specified seed words, we conducted an additional analysis evaluating the performance of SPF when an inappropriate seed word is assigned to a topic. Specifically, we fitted the SPF model with ‘dog’ as a seed word for the ‘Beauty’ topic. The results indicate that SPF effectively recognizes that the term contributes minimal to no informational value in this context. In particular, the inferred variational distribution was $\widetilde{\beta}_{beauty,dog}\sim\Gamma(0.25,5.66)$. These findings underscore SPF’s ability to adaptively assign importance to seed words, thereby reducing the impact of initial specification errors.
4.1.2 Predictive performance
@@
Next, we measure the predictive performance of SPF based on approximate posterior means of the document-topic intensities, where each document vector is a $K$-dimensional vector of approximate mean intensities $\hat{\bm{\theta}}_{d}\in\mathbb{R}_{>0}^{K}$, see Equation 6. According to the Naive Bayes classifier, the topic with the highest approximate mean intensity in $\hat{\bm{\theta}}_{d}$ is assigned as the predicted topic for document $d$. We assess the predictive performance separately for each topic based on precision (correctly assigned to the topic among all assigned to the topic), recall (correctly assigned to the topic among all from the topic) and F1-score (harmonic mean of precision and recall) as well as on aggregate with equal category weights (macro avg) and taking the empirical category imbalance into account (weighted avg). We also provide information on assignment certainty, presenting insights into the model’s confidence in its predictions per category, by determining the fraction of document-topic intensity the topic has the document is assigned to. Specifically, we show the average assignment certainty for true positive (TP) and false positive (FP) predictions.
@@
True label Precision Recall F1-score Assignment certainty Count TP FP Toys 0.92 0.82 0.87 0.68 0.51 8 092 Health 0.75 0.46 0.57 0.64 0.60 6 938 Beauty 0.68 0.79 0.73 0.71 0.54 4 072 Baby 0.71 0.78 0.74 0.70 0.54 4 635 Pets 0.61 0.76 0.74 0.65 0.52 3 792 Grocery 0.51 0.94 0.66 0.78 0.54 2 471 Macro avg 0.71 0.76 0.72 30 000 Weighted avg 0.75 0.73 0.73 30 000
@@
Table 3: Predictive performance of the SPF topic model on Amazon customer feedback, including assignment certainties of TP (true positive) and FP (false positive) predictions. The overall accuracy is 0.73.
The classification performance of the SPF topic model is summarized in Table 3. Clearly, SPF provides excellent classification performance in categorizing Amazon customer feedback across all six product categories, despite slight differences across the categories. The overall accuracy of the model is 0.73, which is consistent with the weighted average F1-score (0.73), accounting for the varying sample sizes across categories. The macro average F1-score (0.72) is slightly lower, reflecting the imbalanced performance among categories.
@@
The model exhibits varying performance across categories. For instance, the highest F1-score is observed for the ‘Toys’ category ($0.87$), reflecting both high precision ($0.92$) and recall ($0.82$). This suggests strong alignment between predicted and true labels. Conversely, the ‘Grocery’ category achieves the lowest F1-score ($0.66$), driven by a significant imbalance between precision ($0.51$) and recall ($0.94$). This discrepancy indicates a tendency to over-predict the ‘Grocery’ category, resulting in higher recall at the cost of precision. The over-prediction of the ‘Grocery’ topic is also reflected in the highest TP assignment certainty score of $0.78$. By contrast, the ‘Health’ category shows an inverted pattern, with a high precision ($0.75$) but a much lower recall ($0.46$), indicating under-representation in predictions. In the ‘Health’ category, the model exhibits the lowest TP assignment certainty at $0.64$, indicating that it is on average less confident in the assignment compared to all other categories when correctly assigning a document. At the same time, the FP assignment certainty is the highest among all categories at $0.60$. This combination of low confidence in true positives and high confidence in false positives highlights the model’s particular struggle with distinguishing health-related feedback, emphasizing the need for further refinement in this category. Table 2 shows that the topic-term intensities of seed words for the ‘Grocery’ and ‘Health’ category are in general not as strong as the ones for the other categories. This likely contributes to the lower classification performance observed for these categories. The correct specification of seed words appears to be a crucial factor in achieving high classification performance, as demonstrated by the results in the ‘Toys’ category. This highlights the model’s particularly strong performance in categories with distinct linguistic characteristics. However, categories like ‘Grocery’ and ‘Health’ reveal areas where the model might benefit from further refinement, such as enhanced domain-specific seed words or adjustments to address label imbalance. Also including an additional unseeded topic which captures feedbacks discussing the purchasing process rather than the product could improve the categorization of the feedback items. The overall accuracy and balanced macro and weighted averages suggest a generally robust model. Further refinement could enhance performance in underperforming categories.
4.2 Comparison to existing methods and scalability
@@
To evaluate the predictive performance and the computational efficiency of the SPF topic model in comparison with other guided topic models, we also fit KeyATM (Eshima et al.,, 2024) and SeededLDA (Watanabe and Zhou,, 2022) to the Amazon corpus. We investigate in particular how the predictive performance and run-times change with the number of documents $D$ in the corpus, varying $D$ from $1k$, over $5k$ and $10k$ to $30k$. The evaluation criteria include the run-time (in minutes) as well as the predictive performance metrics accuracy, precision, recall, and F1-score. These computational experiments were conducted using the hardware setup described in Section 3.3.
@@
For all models, we set the number of topics to $K=6$ and limited the input data to customer feedback only, excluding any additional metadata. Each model was trained using the same set of seed words and the default values for model and prior specifications suggested / implemented in the software packages. Using the same number of MCMC iterations for KeyATM and SeededLDA as well as number of epochs for model fitting regardless of the number of documents $D$ did lead to poor predictive performance results in case only very few documents were included in the corpus. We thus increased the number of MCMC iterations / number of epochs for $D$ equal to $5k$ and $1k$. In particular we used 1 500 MCMC iterations and 150 epochs for $D\in\{10k,30k\}$ and doubled this number to 3 000 MCMC iterations and 300 epochs for $D=5k$ and tripled the number to 4 500 MCMC iterations and 450 epochs for $D=1k$.
@@
Table 4 provides the results for this comparison. The run-time comparison clearly shows that regardless of corpus size SPF always has the shortest run-times compared to SeededLDA and finally KeyATM. The difference in run-time increases with the corpus size. While for a corpus of size $1k$ documents, the run-times of KeyATM and SeededLDA are only approximately 1.5 times the run-time of SPF, the run-times increase by a factor of 3 to 5 times for a corpus of size $30k$ documents.
In terms of accuracy, SPF demonstrates superior or on par performance compared to the other methods across all corpora sizes. For the largest corpus ($D=30k$), SPF achieves an accuracy of $0.73$, equivalent to that of KeyATM and higher than SeededLDA’s $0.65$. As the corpus size decreases, SPF maintains a higher accuracy, notably outperforming the competing methods for smaller corpora. In particular, SPF achieves high accuracy up to a corpus size of $5k$ with a considerable drop in performance only observed in case $1k$ documents are included in the corpus. Also SeededLDA maintains a similar – although at a lower level – accuracy regardless of corpus size with only a slight reduction in case of a small corpus size. KeyATM is most severely effected by a decrease in corpus size with the accuracy values dropping from 0.73 to 0.29. SPF also outperforms the other methods with respect to the other predictive performance criteria such as precision, recall and F1-score.
@@
$30k$ documents $10k$ documents SPF KeyATM SeededLDA SPF KeyATM SeededLDA Time (min:sec) 1:07 5:27 3:35 0:19 1:41 1:04 Accuracy 0.73 0.73 0.65 0.72 0.57 0.63 Precision 0.71 0.73 0.63 0.71 0.72 0.62 Recall 0.76 0.72 0.68 0.75 0.52 0.66 F1-score 0.72 0.72 0.65 0.71 0.55 0.63 $5k$ documents $1k$ documents SPF KeyATM SeededLDA SPF KeyATM SeededLDA Time (min:sec) 0:09 0:44 0:34 0:06 0:09 0:09 Accuracy 0.70 0.44 0.62 0.63 0.29 0.58 Precision 0.68 0.70 0.61 0.61 0.50 0.57 Recall 0.73 0.38 0.64 0.67 0.21 0.59 F1-score 0.69 0.39 0.62 0.62 0.15 0.57
Table 4: SPF vs. existing methods with different corpora sizes.
[…]
@@
4.3 Robustness checks
We systematically vary key parameters of the specification and estimation of the SPF topic model to evaluate how the performance changes under different conditions. Table 5 presents the resulting 13 scenarios and the corresponding predictive performance results obtained when fitting SPF under these scenarios.

### v2

@@ 3.2 Post-processing final inference
The final topic assignment is based on a Naive Bayes classifier [117, see, e.g.,], i.e., the document is assigned to the topic where the per-document posterior mean estimate is maximal.
We employ standard measures used in classification to evaluate the predictive performance of the SPF topic model when used for automatic text classification with known categories. In particular, we use the following metrics separately for each topic: precision (i.e., correctly assigned documents among all documents assigned to the category), recall (i.e., correctly assigned documents among all documents belonging to the category) and F1-score (harmonic mean of precision and recall). In addition, we obtain aggregate measures using either equal category weights (macro avg) or taking the empirical category frequencies into account (weighted avg). We also provide information on assignment certainty, presenting insights into the model’s confidence in its predictions per category, by determining the proportion the document-topic intensity has for the topic the document is assigned to compared to all document-topic intensities. Specifically, we determine the average assignment certainty for true positive (ACTP) and false positive (ACFP) predictions for each category.
Although the primary objective of the SPF topic model is to enhance classification performance, we additionally evaluate topic quality by assessing both topic coherence and topic diversity. Topic coherence is measured using three standard metrics. First, NPMI (Normalized Pointwise Mutual Information; [94]) quantifies the semantic coherence among the top-ranked words of each topic, with higher values indicating better coherence. Second, the UMass score [100] relies on document co-occurrence statistics, where fewer negative values suggest greater topic quality. Third, $C_{V}$ [108] combines several coherence signals using a sliding window, normalized pointwise mutual information, and cosine similarity between word vectors. This metric has been shown to correlate well with human judgment of topic quality. To assess topic diversity, we follow [74] and compute the proportion of unique terms among the top words across all topics, capturing the distinctiveness of the learned topic representations.
@@ 3.3 Computational details
The SPF model is implemented in Python 3.10. It allows Graphics Processing Unit (GPU) support due to its implementation in the TensorFlow environment. The model’s source code is mainly based on TensorFlow 2.18 as well as TensorFlow’s add-on library for probabilistic reasoning, TensorFlow Probability 0.25.0. In its standard implementation, SPF uses a batch size of 1 0241\,024 documents, a learning rate of $0.1$ and trains the model for $150$ epochs. Leveraging TensorFlow’s computational graph and gradient tape functionalities, the implementation enables efficient tracking of operations for automatic differentiation during model training. Tokenization and the construction of the document-term matrix (DTM) are carried out using the Python library scikit-learn [103]. The topic coherence and diversity metrics are computed using the Python library gensim [106], based on the default settings for each metric as provided by the library.
@@ 4 Empirical results
We demonstrate the use of SPF on the Amazon Reviews dataset [86] and four benchmark datasets consisting of text corpora of varying size and text classifications with a varying number of categories. We provide a detailed analysis of the application on the Amazon Reviews dataset to illustrate topic assessment as well as compare SPF to other LDA-based seeded topic models with respect to classification performance, topic coherence and diversity as well as computational efficiency. In addition, we perform a robustness analysis to assess the impact of the hyperparameters. The application to the four benchmark datasets indicate the general applicability of SPF for automatic text classification and the computational efficiency.
When fitting the SPF model in these empirical applications, we set the number of topics to correspond to the number of categories and follow the lines of [116] to construct a balanced lexicon of ten frequently occurring seed words for each category corresponding to a topic. To generate these seed words, we compute the TF-IDF (Term Frequency-Inverse Document Frequency) values for each word within each category. We select the top-10 words from the TF-IDF matrix as the seed words for each topic. This process is conducted in an automatic way to minimize subjectivity and to ensure that the seed word selection process remains as objective as possible.
4.1 Application to the Amazon Reviews dataset
The Amazon Reviews dataset consists of customer feedback on products sold through Amazon from the following six level-1 product categories: health personal care, toys games, beauty, pet supplies, baby products, and grocery gourmet food. Each observation consists of the review text and the information on the product category. To prepare the text data, we apply the following pre-processing steps: text normalization (conversion to lowercase), removal of stop words, and exclusion of words appearing fewer than two times in the corpus. We construct the DTM using a sample of 30 00030\,000 documents, each containing at least seven words. The resulting matrix $\mathbb{Y}$ includes D=30  [^0] D=30\,000 non-empty customer feedback entries and a vocabulary size of V=23 135V=23\,135 unique terms. On average, the documents in $\mathbb{Y}$ contain 42.3 words, with 55% and 9595% quantiles of $15.0$ and $100.0$, respectively. Table 1 presents a summary of the final sample of documents and the pre-specified seed words per product category selected using the proposed automatic procedure.
@@ 4.1 Application to the Amazon Reviews dataset
4.1.1 Topic assessment
We examine the topics inferred by SPF based on the approximate posterior mean topic-term intensities $\hat{\bm{\beta}}_{k}$ for each topic $k$. Table 2 presents the top-14 terms with the highest approximate mean intensities per topic, after removing stop-word-like terms that provide no contextual information. Bold terms represent seed words. The mean term intensity (provided in parentheses) is calculated as the sum of the seeded term intensity and the unseeded term intensity, as defined in Equation (6). These high-intensity words per topic enable the characterization of the topic as well as the assessment of how influential the seed words were.
Clearly the pre-defined seed words exhibit a strong presence among the most pertinent terms for all topics. However, the number of seed words included in the top-14 words with highest intensity varies across topics. For topic ‘Toys‘, all 10 seed words are also included in the list of 14 most pertinent terms for this topic. This number decreases to six for ‘Health’, eight for ‘Beauty’, seven for ‘Baby’ and eight for ‘Pets’. The lowest number of seed words are included in the list of 14 most pertinent terms for the topic ‘Grocery’ where only four out of the ten seed words are listed.
Table 2 reveals that the model effectively prioritizes not only the explicitly defined seed words but also identifies and assigns significant weight to relevant additional terms that are not prespecified as seed words. For example, for topic ‘Health’ the seed word ‘day’ was specified and also ‘days’ is included among the 14 most pertinent terms. For topic ‘Baby’ not only the seed word ‘son’ is included but also ‘daughter’. Inspecting the ‘Baby’ topic further by also assessing additional terms with high intensity indicates that SPF did not only assign high relevance to expected seed words such as ‘seat’ (14.88) and ‘son’ (12.07), but also recognized terms like ‘bed’ (6.71) and ‘sleep’ (6.31) as highly pertinent to the topic. These terms align with the product subcategory ‘sleep positioners’, which falls under the broader ‘Baby’ category, demonstrating the model’s nuanced understanding of topic content and its ability to discern contextually important terms.
The topic ‘Grocery’ fails to capture most of the seed words among the 14 most pertinent terms. However, the terms with high intensity suggest that this topic captured an additional aspect in customer reviews which relates not to the product but to the purchase process. For example, terms like ‘store’ ($7.34$) and ‘shipping’ ($7.20$) have a high prevalence in the ‘Grocery’ topic.
@@ 4.1.1 Topic assessment
Table 2 also illustrates that SPF is able to assign distinctly different intensities to the seed words as well as other terms with high intensity within their respective topics. E.g., within the ‘Pets’ category, ‘dog’ ($24.57$) plays a dominant role, whereas other seed words, like ‘food’ ($7.41$), display a markedly lower mean intensity. This contrast highlights the ability of SPF to estimate the uneven influence of seed words in defining a topic. This property of the SPF topic model is important to also mitigate the risk of potential misspecifications that may arise due to incomplete domain knowledge during the initial selection of seed words. To empirically assess the influence of misspecified seed words, we conducted an additional analysis evaluating the performance of SPF when an inappropriate seed word is assigned to a topic. Specifically, we fitted the SPF model with ‘dog’ as a seed word for the ‘Beauty’ topic. The results indicate that SPF effectively recognizes that the term contributes minimal to no informational value in this context. In particular, the inferred variational distribution was $\widetilde{\beta}_{beauty,dog}\sim\Gamma(0.25,5.66)$. These findings underscore SPF’s ability to adaptively assign importance to seed words, thereby reducing the impact of initial specification errors.
4.1.2 Classification performance
@@
Next, we measure the classification performance of SPF based on approximate posterior means of the document-topic intensities, where each document vector is a $K$-dimensional vector of approximate mean intensities $\hat{\bm{\theta}}_{d}\in\mathbb{R}_{>0}^{K}$, see Equation (5). According to the Naive Bayes classifier, the topic with the highest approximate mean intensity in $\hat{\bm{\theta}}_{d}$ is assigned as the predicted topic for document $d$. We assess the classification performance separately for each category.
@@
Category Precision Recall F1-score ACTP ACFP Toys 0.92 0.82 0.87 0.68 0.51 Health 0.75 0.46 0.57 0.64 0.60 Beauty 0.68 0.79 0.73 0.71 0.54 Baby 0.71 0.78 0.74 0.70 0.54 Pets 0.61 0.76 0.74 0.65 0.52 Grocery 0.51 0.94 0.66 0.78 0.54 Macro avg 0.71 0.76 0.72 Weighted avg 0.75 0.73 0.73
@@
Table 3: Classification performance of the SPF topic model on Amazon customer feedback, including assignment certainty of true positives (ACTP) and false positives (ACFP). The overall accuracy is 0.73.
@@
The classification performance of the SPF topic model is summarized in Table 3. Clearly, SPF provides excellent classification performance in categorizing Amazon customer feedback across all six product categories, despite slight differences among categories. The overall accuracy of the model is 0.73, which is consistent with the weighted average F1-score (0.73), accounting for the varying sample sizes across categories. The macro average F1-score (0.72) is slightly lower, reflecting the imbalanced performance among categories. For instance, the highest F1-score is observed for the ‘Toys’ category ($0.87$), reflecting both high precision ($0.92$) and recall ($0.82$). This suggests a strong alignment between the predicted and true labels. Conversely, the ‘Grocery’ category achieves the lowest F1-score ($0.66$), driven by a significant imbalance between precision ($0.51$) and recall ($0.94$). This discrepancy indicates a tendency to over-predict the ‘Grocery’ category, resulting in higher recall at the cost of precision. The over-prediction of the ‘Grocery’ topic is also reflected in the highest ACTP score of $0.78$. By contrast, the ‘Health’ category shows an inverted pattern, with a high precision ($0.75$) but a much lower recall ($0.46$), indicating under-representation in predictions. In the ‘Health’ category, the model exhibits the lowest ACTP score at $0.64$, indicating that SPF is on average less confident in the assignment compared to all other categories when correctly assigning a review. At the same time, the ACFP score is the highest among all categories at $0.60$. This combination of low confidence in true positives and high confidence in false positives highlights the model’s particular struggle with distinguishing health-related feedback, emphasizing the need for further refinement in this category. Table 2 shows that the topic-term intensities of seed words for the ‘Grocery’ and ‘Health’ category are in general not as strong as the ones for the other categories. This likely contributes to the lower classification performance observed for these categories. The correct specification of seed words appears to be a crucial factor in achieving high classification performance, as demonstrated by the results in the ‘Toys’ category. This highlights the model’s particularly strong performance in categories with distinct linguistic characteristics. However, categories like ‘Grocery’ and ‘Health’ reveal areas where the model might benefit from further refinement, such as enhanced domain-specific seed words or adjustments to address label imbalance. In addition, including an additional unseeded topic that captures feedbacks discussing the purchasing process rather than the product could improve the categorization of the feedback items. Nevertheless, the overall accuracy and balanced macro and weighted averages suggest a generally robust model, even if further refinement could enhance performance in underperforming categories.
4.1.3 Comparison to existing methods and scalability
@@
To evaluate the classification performance and the computational efficiency of the SPF topic model in comparison with other guided topic models, we also fit KeyATM [76] and SeededLDA [116] to the Amazon corpus. We complement this with a comparison to the standard PF model. We investigate in particular how the classification performance and run-times change with the number of documents $D$ in the corpus, varying $D$ from $1k$, over $5k$ and $10k$ to $30k$. The evaluation criteria include the run-time (in minutes) as well as the classification performance metrics accuracy, precision, recall, and F1-score. These computational experiments were conducted using the hardware setup described in Section 3.3.
When fitting the standard PF topic model, the inferred topics are unlabeled and their ordering is arbitrary. Evaluating the classification accuracy requires identifying an alignment between the predicted topics and the true category labels. To do so, we apply the Hungarian algorithm [91] to obtain the permutation of predicted topics that maximizes accuracy. This approach is a form of optimal label permutation commonly used in clustering and unsupervised learning evaluation, and is equivalent to solving the linear sum assignment problem [101]. Based on this optimal mapping, the predicted labels are accordingly remapped and the classification metrics calculated.
@@
For all models, we set the number of topics to $K=6$ and limited the input data to customer feedback only, excluding any additional metadata. Each model was trained using the same set of seed words and the default values for model and prior specifications suggested / implemented in the software packages. Using the same number of MCMC iterations for KeyATM and SeededLDA as well as the number of epochs for model fitting regardless of the number of documents $D$ led to poor predictive performance results in the case where only very few documents were included in the corpus. We thus increased the number of MCMC iterations / number of epochs for $D$ equal to $5k$ and $1k$. In particular, we used 1 500 MCMC iterations and 150 epochs for $D\in\{10k,30k\}$ and doubled this number to 3 000 MCMC iterations and 300 epochs for $D=5k$ and tripled the number to 4 500 MCMC iterations and 450 epochs for $D=1k$.
@@
Table 4 provides the results for this comparison. The run-time comparison clearly shows that regardless of corpus size, SPF has a comparable run-time to PF and it always has the shortest run-times compared to both SeededLDA and KeyATM. The difference in run-time increases with the corpus size. While for a corpus of size $1k$ documents, the run-times of KeyATM and SeededLDA are only approximately 1.5 times the run-time of SPF, the run-times increase by a factor of 3 to 5 times for a corpus of size $30k$ documents.
@@
$30k$ documents $10k$ documents SPF KeyATM SeededLDA PF SPF KeyATM SeededLDA PF Time (min:sec) 1:07 5:27 3:35 1:04 0:19 1:41 1:04 0:19 Accuracy 0.73 0.73 0.65 0.55 0.72 0.57 0.63 0.50 Precision 0.71 0.73 0.63 0.52 0.71 0.72 0.62 0.47 Recall 0.76 0.72 0.68 0.56 0.75 0.52 0.66 0.51 F1-score 0.72 0.72 0.65 0.53 0.71 0.55 0.63 0.48 $5k$ documents $1k$ documents SPF KeyATM SeededLDA PF SPF KeyATM SeededLDA PF Time (min:sec) 0:09 0:44 0:34 0:09 0:06 0:09 0:09 0:07 Accuracy 0.70 0.44 0.62 0.61 0.63 0.29 0.58 0.32 Precision 0.68 0.70 0.61 0.59 0.61 0.50 0.57 0.33 Recall 0.73 0.38 0.64 0.64 0.67 0.21 0.59 0.35 F1-score 0.69 0.39 0.62 0.60 0.62 0.15 0.57 0.32
Table 4: Classification performance across models and corpus sizes. Bold: The highest score. Underline: The second highest score.
In terms of accuracy, SPF demonstrates superior or on par performance compared to the other methods across all corpora sizes. The unguided PF topic model generally exhibits the weakest predictive performance among the evaluated methods. This is expected due to its completely unsupervised nature, which lacks any form of guidance during training. In contrast, the SPF model – despite incorporating only a minimal amount of domain knowledge – achieves substantially better predictive results, highlighting the effectiveness of even weak supervision in improving classification performance with topic models when categories can be pre-specified. For the largest corpus ($D=30k$), SPF achieves an accuracy of $0.73$, equivalent to that of KeyATM and higher than SeededLDA’s $0.65$, while PF only has an accuracy of 0.55. As corpus size decreases, SPF maintains a higher accuracy, notably outperforming competing methods for smaller corpora. In particular, SPF achieves high accuracy up to a corpus size of $5k$ with a considerable drop in performance only observed in case $1k$ documents are included in the corpus. Also, SeededLDA maintains a similar – although at a lower level – accuracy regardless of corpus size with only a slight reduction in the case of a small corpus size. KeyATM is most severely affected by a decrease in corpus size with the accuracy values dropping from 0.73 to 0.29. SPF also outperforms the other methods with respect to the other predictive performance criteria such as precision, recall and F1-score.
Table 5 compares SPF, KeyATM, SeededLDA, and plain PF in terms of topic coherence and diversity across varying corpus sizes. For each model, the scores are computed based on the top-10 highest-ranked terms per topic. Overall, SPF exhibits competitive performance in the topic coherence metrics. While SeededLDA achieves the highest $C_{V}$ coherence scores in most settings, SPF consistently matches or outperforms its competitors on NPMI and UMass. For example, at corpus sizes of 30kk and 10kk, SPF achieves UMass scores of 1.80-1.80 and 1.75-1.75, respectively – higher (i.e., better) than those of SeededLDA, and comparable to or slightly below those of KeyATM. At the smallest data size (1kk documents), SPF clearly outperforms both baselines on UMass, highlighting its robustness in low-resource scenarios. Interestingly, SPF and PF yield similar results in coherence and diversity across all corpus sizes.
In terms of topic diversity, SPF generally falls between KeyATM and SeededLDA, while performing comparably to PF. It provides more diverse topics than KeyATM on the 30kk dataset but remains below SeededLDA, which achieves the highest diversity scores across most corpus sizes. At smaller scales (i.e., 5kk and 1kk documents), KeyATM surpasses SPF in diversity, although this comes at the cost of lower coherence. These results highlight the design trade-offs between models. We emphasize that while the primary objective of the SPF model is to enhance classification performance, it adeptly balances coherence and diversity, with a consistent advantage in maintaining topic quality under data constraints. Tables 4 and 5 together highlight the importance of not relying solely on topic quality metrics when evaluating models for classification tasks. For instance, in the 5kk document setting, KeyATM attains a higher topic diversity score ($0.82$) than SPF ($0.68$). Nevertheless, this increased diversity is accompanied by a markedly lower predictive accuracy – $0.44$ for KeyATM compared to $0.70$ for SPF – illustrating that greater topic diversity does not necessarily imply superior classification performance.
$30k$ documents $10k$ documents SPF KeyATM SeededLDA PF SPF KeyATM SeededLDA PF NPMI 0.15 0.16 0.14 0.15 0.11 0.12 0.11 0.10 UMass -1.80 -1.68 -2.03 -1.78 -1.75 -2.02 -2.22 -1.74 $C_{V}$ 0.58 0.59 0.64 0.57 0.58 0.61 0.64 0.56 Diversity 0.67 0.57 0.83 0.67 0.63 0.73 0.83 0.63 $5k$ documents $1k$ documents SPF KeyATM SeededLDA PF SPF KeyATM SeededLDA PF NPMI 0.13 0.11 0.08 0.14 0.07 0.02 0.02 0.11 UMass -1.81 -2.22 -2.57 -1.88 -1.71 -4.63 -3.36 -1.89 $C_{V}$ 0.54 0.57 0.59 0.52 0.54 0.52 0.52 0.51 Diversity 0.68 0.82 0.87 0.68 0.63 0.97 0.92 0.67
Table 5: Coherence and diversity comparison across different corpora sizes. Bold: The highest score. Underline: The second highest score.
To evaluate the scalability of SPF, we conduct a bootstrap experiment where we draw documents with replacement from the available corpus to obtain corpora of different size. In particular, we increase the number of documents $D$ in increments of $100k$, up to a total of D=1 000  [^0] D=1\,000\,000 documents. This bootstrap experiment was conducted exclusively with the SPF model, as both KeyATM and SeededLDA were unable to handle such large corpora on the hardware described in Section 3.3. Figure 3 visualizes the run-times observed, indicating a roughly linear increase in run-times as the corpus size grows. This increase may be attributed to the increase in the model’s local variational parameters $\phi^{\mathsf{rte}}_{\theta}$ and $\phi^{\mathsf{shp}}_{\theta}$ with the number of documents. Figure 3 shows that SPF successfully processes the corpus with 1 000 0001\,000\,000 documents in approximately 2 hours, demonstrating its ability to handle large-scale datasets efficiently.
@@
4.1.4 Robustness checks
We systematically vary key parameters of the specification and estimation of the SPF topic model to evaluate how the performance changes under different conditions. Table 6 presents the resulting 13 scenarios and the corresponding classification performance results obtained when fitting SPF in these scenarios.